# Guided simulation of conditioned chemical reaction networks

### v2

@@ 6.2.1. Death model and comparison to Golightly & Sherlock, (2019)
It is a well-known result that $X(T)\sim\mathrm{Binom}\left(x_{0},e^{-cT}\right)$, so we will use this density for comparison. In the following experiment, we use (6.4) to estimate the mass function of the $\mathrm{Binom}\left(x_{0},e^{-cT}\right)$-distribution with $x_{0}=50$, $T=1$ and $c=1/2$. We consider four choices for $g$:
@@
  • $g$ chosen using the density of a reversed (decreasing) constant rate Poisson process with intensity $\theta$. We took $\theta=cv$ (this choice ensures equivalence by Theorem 4.2 of Corstanje et al., (2023)).
@@
Figure 4. The percentages of sample paths with $X(T)=v$ for each $v$ for the three methods with the same settings as Figure 3
@@ Appendix B Proofs for Section 5
    $\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}^{h}_{t}\left[\frac{g(t,X_{t})}{h(t,X_{t})}\mathbf{1}_{A_{j}[X^{t}]}\right]=\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A[X^{t_{k}}]}\right]}{h(0,x_{0})},$
@@ Appendix D Change of generator under change of measure
$\frac{\,\mathrm{d}\tilde{\mathbb{P}}_{t}}{\,\mathrm{d}{\mathbb{P}}_{t}}=E^{g}(t):=\frac{g(t,X(t))}{g(0,x_{0})}\exp\left(-\int_{0}^{t}\frac{\mathcal{A}g}{g}(s,X_{s})\,\mathrm{d}s\right),$
@@
$\tilde{D}^{f}(t):=f(t,X_{t})-\int_{0}^{t}(\tilde{\mathcal{A}}f)(s,X_{s})\,\mathrm{d}s$

### v3

@@ 6.2.1. Death model and comparison to Golightly & Sherlock, (2019)
It is a well-known result that $X(T)\sim\mathrm{Binom}\left(x_{0},e^{-cT}\right)$, so we will use this mass function for comparison. In the following experiment, we use (6.4) to estimate the mass function of the $\mathrm{Binom}\left(x_{0},e^{-cT}\right)$-distribution with $x_{0}=50$, $T=1$ and $c=1/2$. We consider four choices for $g$:
@@
  • $g$ chosen using the density of a reversed (decreasing) constant rate Poisson process with rate constant $\theta$. We took $\theta=cv$ (this choice ensures equivalence by Theorem 4.2 of Corstanje et al., (2023)).
@@
Figure 4. The percentage of paths ending in the point of conditioning ($v$) versus $v$ for the four methods considered (with the same settings as in Figure 3).
@@ Appendix B Proofs for Section 5
    $\lim_{j\to\infty}\lim_{t\uparrow t_{k}}{\mathbb{E}}^{h}_{t}\left[\frac{g(t,X(t))}{h(t,X(t))}\mathbf{1}_{A_{j}[X^{t}]}\right]=\frac{{\mathbb{E}}\left[g(t_{k},X(t_{k}))\mathbf{1}\{L_{k}X(t_{k})=v_{k}\}\mathbf{1}_{A[X^{t_{k}}]}\right]}{h(0,x_{0})},$
@@ Appendix D Change of generator under change of measure
$\frac{\,\mathrm{d}\tilde{\mathbb{P}}_{t}}{\,\mathrm{d}{\mathbb{P}}_{t}}=E^{g}(t):=\frac{g(t,X(t))}{g(0,x_{0})}\exp\left(-\int_{0}^{t}\frac{\mathcal{A}g}{g}(s,X(s))\,\mathrm{d}s\right),$
@@
$\tilde{D}^{f}(t):=f(t,X(t))-\int_{0}^{t}(\tilde{\mathcal{A}}f)(s,X(s))\,\mathrm{d}s$